11)   Find the proceeds of a 5 year non-interest bearing note for $6500 discounted 2.5 years before maturity if money earns 6% p.a. compounded quarterly.
Answer #1

I = 6% / 4 = 1.5% per quarter

t = 2.5 yrs = 2.5 * 4 = 10 quarters

P = F / (1+I)^t = 6500 / (1+0.015)^10

= 6500 / (1.015)^10

= 5600.83